# Simulating and Understanding Deceptive Behaviors in Long-Horizon Interactions

### v2

@@ 3.2 Supervisor Agent
The supervisor state is initialized at a neutral point $\mathbf{s}_{1,0}=(0,0,0)$, reflecting baseline trust, satisfaction, and comfort before any interaction. Here (i) $s^{\text{trust}}_{i,j}$ measures perceived integrity and reliability of the performer, grounded in trust calibration literature in supervisory relationships (Glikson & Woolley, 2020); (ii) $s^{\text{satisfaction}}_{i,j}$ quantifies task-focused quality, reflecting alignment between the performer’s outputs and the requirements of $T_{i}$; and (iii) $s^{\text{comfort}}_{i,j}$ encodes relational comfort level in the collaboration  (Jarrahi, 2018).
Memory. The supervisor agent maintains a hybrid memory system $\mathcal{M}$ that retains both short-term detail and long-term summaries, similar to designs in prior work (Zhong et al., 2024; Park et al., 2023a; Zhang et al., 2025). Recent $K$ interactions are stored verbatim, while earlier exchanges are compressed into abstractive summaries. Artifacts from the task stream, such as documents or structured outputs (e.g., .txt or .json files), are also preserved. This memory structure enables the supervisor to evaluate not only local task performance but also global coherence across tasks, making it sensitive to inconsistencies, omissions, or fabrications that may unfold over time.
@@ 3.3 Deception Auditor
Auditor inputs and context. The deception auditor is executed after the completion of the full long-horizon interaction. For each interaction step $(i,j)$, the auditor receives as input (i) a fixed rubric of deception indicators (Appendix A.5); (ii) the current interaction slices $\mathcal{H}_{i,j}=(T_{i},e_{i},a_{i,j},f_{i,j})$ corresponding to each task $T_{i}$ and its associated rounds; and (iii) the complete trajectory $\mathcal{H}$ accumulated over the entire interactions. By conditioning its judgments on $\mathcal{H}_{i,j}$ in the context of the global history $\mathcal{H}$, the auditor evaluates not only local behaviors but also longitudinal coherence across tasks. This is motivated by meta-analytic evidence showing that human lie–truth judgments are only marginally above chance when assessed turn by turn, whereas many deceptive strategies become identifiable only when sequences are examined in aggregate (BondJr. & DePaulo, 2006).
@@ 4.1 Experimental Setup
We evaluate our framework on a diverse set of frontier language models, including closed-source models such as GPT-4o (OpenAI, 2024), GPT-o3 and o4-mini (OpenAI, 2025c), Gemini 2.5 Pro (Google, 2025), Claude Sonnet 4 (Anthropic, 2025a), Claude Opus 4.1 (Anthropic, 2025b), and Grok 4 (xAI, 2025), as well as open-source releases such as GPT-OSS-120B (OpenAI, 2025b), Qwen3 (Yang et al., 2025), DeepSeek-V3-0324 (DeepSeek, 2025b), DeepSeek-R1-0528 (DeepSeek, 2025a), and DeepSeek-V3.1 (DeepSeek, 2025c). We used the default API and inference parameters without modification, with the context length set to the maximum supported by each model. We set the maximum number of rounds per task $R_{\text{max}}=3$. In the main experiments, each model was run for 20 independent trials under the same random seed, ensuring that event sequences were identical across models for fair comparison. We consistently use 14 tasks for all models, defined in the Appendix A.1. Additional experimental details and results on a different domain are provided in the Appendix B.
@@ 4.2 Results
How does deception rate vary across LLMs? We compare deception rates across models, defined as the proportion of interactions flagged by the auditor as deceptive, among all the interactions in a trajectory. Table 1 summarizes the results of the average deception rate across 20 runs, where each run represents a full long-horizon trajectory. Models separate into clear tiers. Claude Sonnet-4 remains the least deceptive with an average deception rate of $0.214$, indicating sporadic deception. In contrast, models including GPT-4o, DeepSeek R1-0528, and gpt-oss-120b show markedly higher rates (0.6370.6370.7610.761), suggesting a greater tendency to deceive under pressure. Finally, the highest deception rate appears in DeepSeek V3-0324, which is $0.793$. These findings demonstrate that deception is not an inevitable byproduct of long-horizon interaction but rather varies systematically across architectures, with certain model families consistently more prone to deceptive strategies.
@@
What type of deceptions do models employ? We next analyze the distribution of deception strategies, as classified by the auditor under Interpersonal Deception Theory (IDT). Figure 3 summarizes the relative proportions of falsification (green), concealment (yellow), and equivocation (blue) across models. Across the board, falsification dominates as the primary strategy: most models rely heavily on fabricating information. This tendency is especially pronounced in Gemini 2.5 Pro and gpt-oss-120b, where falsification accounts for the vast majority of deceptive responses. In contrast, models such as GPT-4o-1120 and DeepSeek V3-0324, while still led by falsification, allocate relatively larger shares to concealment and equivocation, signaling a broader mix of evasive tactics. These findings suggest that although falsification is the default strategy across systems, models differ in how much they diversify into subtler forms of deception, reflecting differences in inductive biases and training regimes.
Figure 4: Relationship between deception rate ($y$-axis) and supervisor agent’s states: trust (left), satisfaction (middle), and relational comfort (right).
How does deception affect the supervisor agent’s trust? In our framework, an interesting aspect is to examine the relationship between deception and the supervisor agent’s trust level. Recall that at the end of each task completion, the supervisor agent updates a trust score, represented by the state variable $s^{\text{trust}}\in[-1,1]$, which captures the perceived integrity and reliability of the performer. After all 14 tasks in the stream are completed, these scores are averaged to obtain a trajectory-level measure of trust. We then compare this trajectory-level trust score with the deception rate observed for the same trajectory. The results in Figure 4 show a strong anti-correlation: models with higher deception rates systematically achieve lower trust. For instance, Gemini 2.5 Pro maintains a relatively low deception rate and correspondingly high trust, whereas DeepSeek V3-0324 displays both elevated deception rates ($\approx 0.8$) and sharply negative trust scores (<0.75<-0.75). This affirms that deception erodes trust in the performer-supervisor interaction. The figure also shows that models in the same tier have a similar trust score. Similarly, we provide analysis on how deception affects the supervisor agent’s satisfaction score and relational comfort score in Figure 4, which reveals the same trend of anti-correlation and the tendency of trickier deception behaviors in frontier models.
@@ 5 Discussion
Figure 5: Impact of events on deceptive behaviors. Left: Event category vs. deception type. Right: Pressure level vs. deception rate.
@@ 5.1 Control Study
How does the pressure level of events impact deception? We further examine the role of event pressure by conditioning deception rates on the pressure level at which events are injected. Recall that each category contains events spanning four levels of pressure, from low to critical. Figure 5 (right) compares Deepseek V3.1 and GPT-4o under varying levels of pressure. Results reveal that for both models, deception rates rise steadily as pressure increases from low to high, suggesting that event pressure acts as an amplifier of deceptive strategies for LLMs. While surprisingly, for GPT-4o, the deception rates drop at the level of critical pressure. Through qualitative analysis (see Appendix C.2 for more details), we found that under critical pressure, GPT-4o tends to explicitly respond to the event with a safety-aware statement, encouraging itself to be honest and reliable. Such behavior does not show when exposed in events with high or medium pressure.
@@ 5.2 Qualitative Study
We conduct qualitative case studies on deception behaviors in long-horizon tasks, with full details in Appendix C. Our analysis reveals a recurring issue of “chain of deception”: the performer agent gradually deviates from its constraints, and—unless explicitly corrected by the supervisor agent—eventually fabricates information. This behavior begins subtly but quickly escalates over multiple steps. We also observe that the feedback provided by the supervisor agent is frequently ignored by the performer agent, leading to repeated instances of deceptive behavior. These findings underscore the risks associated with deploying LLMs in long-horizon or loosely supervised scenarios, particularly in tasks that demand sustained alignment over time.
@@ 6 Conclusion
We introduced the first long-horizon simulation framework for probing deception in large language models, integrating structured task streams, probabilistic event systems, and performer–supervisor interactions with independent auditing. Our experiments across 11 frontier models reveal that deception is not uniformly distributed, increases event pressure, and is strongly anti-correlated with supervisor trust. These findings highlight deception as an emergent phenomenon in long-horizon interactions, overlooked by short-form benchmarks, and suggest that training regimes and inductive biases shape deceptive tendencies. By grounding our framework in social science insights and systematically quantifying deceptive behavior, we provide both a methodological foundation and empirical evidence to guide the design of more trustworthy LLMs in sustained, high-stakes settings.
@@
This work advances the evaluation of large language models by introducing the first framework for systematically studying deception in long-horizon interactions. While our findings provide valuable insights into how deceptive strategies emerge, they also highlight potential risks. Deception is a socially consequential behavior: when deployed in trust-sensitive settings such as education, healthcare, or enterprise decision-making, models that obscure errors or fabricate evidence could undermine user trust and cause real-world harm. By quantifying deception across both closed- and open-source frontier models, our study provides guidance for the development of safer systems and more robust evaluation protocols. At the same time, we acknowledge that exposing deceptive tendencies might inform adversarial uses, and we stress that our framework should be applied responsibly, with the goal of improving transparency and alignment rather than enabling misuse.
@@
To ensure the reproducibility of our research, we provide comprehensive details of our framework and experimental setup. The full implementation of our simulation framework, including the task streams, the theoretically grounded event system, and the complete prompts for the Supervisor Agent and Deception Auditor, is detailed in Appendix A.1, A.2, A.4, and A.5. Details of our experimental setup, including the evaluated models, key parameters, and the use of a seed to ensure identical event sequences for all trials, are described in Section 4.1 and Appendix B.1. Our code repository is available at https://github.com/deeplearning-wisc/LongHorizonDeception.
@@ A.2 Event System
Each event belongs to one of the five categories: (1) Goal Conflict, (2) Competitive Interaction, (3) Moral Dilemma, (4) Authority Directive, and (5) Information Gap. These five categories subsume established stressor and ethics taxonomies in organizational behavior and management science, covering the major axes of exogenous pressure identified in prior work. Specifically, Kahn et al. (1964) identified role conflict and role ambiguity, directly mapping to Goal Conflict and Information Gap; Cooper & Marshall (1976) classified job-related stressors into job factors, role in organization, relationships, and organizational structure, aligning with Goal Conflict, Competitive Interaction, and Authority Directive; Podsakoff et al. (2007) empirically distinguished conflict, workload, job insecurity, and organizational politics, mapping to Moral Dilemma, Competitive Interaction, and Authority Directive. In addition, behavioral ethics reviews and meta-analyses, such as Treviño et al. (2006) and Kish-Gephart et al. (2010), highlighted moral awareness and value trade-offs under uncertainty, anchoring Moral Dilemma. The definition of each category, as well as the supporting literature and examples, is provided below:
@@
In addition to categorical mechanisms, prior research shows that the magnitude of consequences and temporal immediacy of an issue systematically alter ethical and strategic choices (Jones, 1991). Accountability research further demonstrates that stronger evaluative scrutiny amplifies distortion risks under pressure (Lerner & Tetlock, 1999), while decision-making studies show that time pressure itself degrades judgment quality and shifts strategy use (Svenson & Maule, 1993). These findings establish pressure level as a necessary dimension, orthogonal to mechanism, for modeling how the same event scales from low-stakes to critical-stakes contexts.

### v3

@@ 3.2 Supervisor Agent
The supervisor state is initialized at a neutral point $\mathbf{s}_{1,0}=(0,0,0)$, reflecting baseline trust, satisfaction, and comfort before any interaction. Here (i) $s^{\text{trust}}_{i,j}$ measures perceived integrity and reliability of the performer, grounded in trust calibration literature in supervisory relationships (Glikson and Woolley, 2020); (ii) $s^{\text{satisfaction}}_{i,j}$ quantifies task-focused quality, reflecting alignment between the performer’s outputs and the requirements of $T_{i}$; and (iii) $s^{\text{comfort}}_{i,j}$ encodes relational comfort level in the collaboration (Jarrahi, 2018). For further discussion on these distinct states, see Appendix A.4.
Memory. The supervisor agent maintains a hybrid memory system $\mathcal{M}$ that retains both short-term detail and long-term summaries, similar to designs in prior work (Zhong et al., 2024; Park et al., 2023a; Zhang et al., 2025b). Recent $K$ interactions are stored verbatim, while earlier exchanges are compressed into abstractive summaries. Artifacts from the task stream, such as documents or structured outputs (e.g., .txt or .json files), are also preserved. This memory structure enables the supervisor to evaluate not only local task performance but also global coherence across tasks, making it sensitive to inconsistencies, omissions, or fabrications that may unfold over time.
@@ 3.3 Deception Auditor
Auditor inputs and context. The deception auditor is executed after the completion of the full long-horizon interaction. For each interaction step $(i,j)$, the auditor receives as input (i) a fixed rubric of deception indicators (Appendix A.5); (ii) the current interaction slices $\mathcal{H}_{i,j}=(T_{i},e_{i},a_{i,j},f_{i,j})$ corresponding to each task $T_{i}$ and its associated rounds; and (iii) the complete trajectory $\mathcal{H}$ accumulated over the entire interactions. By conditioning its judgments on $\mathcal{H}_{i,j}$ in the context of the global history $\mathcal{H}$, the auditor evaluates not only local behaviors but also longitudinal coherence across tasks. This is motivated by meta-analytic evidence showing that human lie–truth judgments are only marginally above chance when assessed turn by turn, whereas many deceptive strategies become identifiable only when sequences are examined in aggregate (BondJr. and DePaulo, 2006).
@@ 4.1 Experimental Setup
We evaluate LH-Deception on a diverse set of frontier language models, including closed-source models such as GPT-4o (OpenAI, 2024), GPT-o3 and o4-mini (OpenAI, 2025c), Gemini 2.5 Pro (Google, 2025), Claude Sonnet 4 (Anthropic, 2025b), Claude Opus 4.1 (Anthropic, 2025a), and Grok 4 (xAI, 2025), as well as open-source releases such as GPT-OSS-120B (OpenAI, 2025b), Qwen3 (Yang et al., 2025), DeepSeek-V3-0324 (DeepSeek, 2025c), DeepSeek-R1-0528 (DeepSeek, 2025a), and DeepSeek-V3.1 (DeepSeek, 2025b). We used the default API and inference parameters without modification, with the context length set to the maximum supported by each model. We set the maximum number of rounds per task $R_{\text{max}}=3$. In the main experiments, each model was run for 20 independent trials under the same random seed, ensuring that event sequences were identical across models for fair comparison. We consistently use 14 tasks for all models, defined in the Appendix A.1. Additional experimental details and results on a different domain are provided in the Appendix B.1. We provide human evaluations to verify the reliability of the LLM auditor in Appendix B.3.
@@ 4.2 Results
How does deception rate vary across LLMs? We compare deception rates across models, defined as the proportion of interactions flagged by the auditor as deceptive, among all the interactions in a trajectory. Table 1 summarizes the results of the average deception rate across 20 runs, where each run represents a full long-horizon trajectory. Models separate into clear tiers. Claude Sonnet-4 remains the least deceptive with an average deception rate of $0.214$, indicating sporadic deception. In contrast, models including GPT-4o, DeepSeek R1-0528, and gpt-oss-120b show markedly higher rates (0.6370.6370.7610.761), suggesting a greater tendency to deceive under pressure. Finally, the highest deception rate appears in DeepSeek V3-0324, which is $0.793$. These findings demonstrate that deception is not an inevitable byproduct of long-horizon interaction but rather varies across models, with certain model families consistently more prone to deceptive strategies.
@@
What type of deceptions do models employ? We next analyze the distribution of deception strategies, as classified by the auditor under Interpersonal Deception Theory (IDT). Figure 3 summarizes the relative proportions of falsification (green), concealment (yellow), and equivocation (blue) across models. Across the board, falsification dominates as the primary strategy: most models rely heavily on fabricating information. This tendency is especially pronounced in Gemini 2.5 Pro and gpt-oss-120b, where falsification accounts for the vast majority of deceptive responses. In contrast, models such as GPT-4o-1120 and DeepSeek V3-0324, while still led by falsification, allocate relatively larger shares to concealment and equivocation, signaling a broader mix of evasive tactics. These behavioral differences reflect variations in their underlying training objectives and inductive biases, leading to distinct strategies: some models rely on pure fabrication while others diversify into subtler evasive tactics.
Figure 4: Relationship between deception rate ($y$-axis) and supervisor agent’s states: trust (left), satisfaction (middle), and relational comfort (right). Full data with standard deviation is available in Appendix 2.
How does deception affect the supervisor agent’s trust? In our framework, an interesting aspect is to examine the relationship between deception and the supervisor agent’s trust level. Recall that at the end of each task completion, the supervisor agent updates a trust score, represented by the state variable $s^{\text{trust}}\in[-1,1]$, which captures the perceived integrity and reliability of the performer. After all 14 tasks in the stream are completed, these scores are averaged to obtain a trajectory-level measure of trust. We then compare this trajectory-level trust score with the deception rate observed for the same trajectory. The results in Figure 4 show a strong anti-correlation: models with higher deception rates systematically achieve lower trust. For instance, Gemini 2.5 Pro maintains a relatively low deception rate and correspondingly high trust, whereas DeepSeek V3-0324 displays both elevated deception rates ($\approx 0.8$) and sharply negative trust scores (<0.75<-0.75). This affirms that deception erodes trust in the performer-supervisor interaction. Similarly, we provide analysis on how deception affects the supervisor agent’s satisfaction score and relational comfort score in Figure 4, which reveals the same trend of anti-correlation and the tendency of trickier deception behaviors in frontier models. While the three scores are correlated, they capture distinct aspects of the relational cost, see Appendix A.4 for a qualitative example.
@@ 5 Discussion
5.1 Comparison with Existing Benchmarks
Our framework measures a different construct than static, single-turn benchmarks. Static benchmarks (e.g., Browne, 2025; Huang et al., 2025) measure capacity for deception or failure in a single instance, whereas our framework measures emergent deception under sustained, relational pressure. This gap is evidenced by the results. For example, GPT-4o, which shows a 29.3% deception rate on DeceptionBench (Huang et al., 2025), exhibits a significantly higher deception rate in our more complex, sustained environment (63.7%). The contrast also exists when compared to SnitchBench (Browne, 2025), which measures single-turn refusal: o4-mini appear 5.0% failure rate in static, single-turn refusal tasks; However, when placed in our long-horizon framework, it demonstrates much higher deception rates (31.3%). This suggests that a model can pass a static test and still fail catastrophically when deployed in a dynamic, long-horizon setting.
@@ 5.2 Control Study
Figure 5: Impact of events on deceptive behaviors. Left: Event category vs. deception type. Right: Pressure level vs. deception rate. Full data with standard error is available in Appendix 2.
How does the pressure level of events impact deception? We further examine the role of event pressure by conditioning deception rates on the pressure level at which events are injected. Recall that each category contains events spanning four levels of pressure, from low to critical. Figure 5 (right) compares Deepseek V3.1 and GPT-4o under varying levels of pressure. Results reveal that for both models, deception rates rise steadily as pressure increases from low to high, suggesting that event pressure acts as an amplifier of deceptive strategies for LLMs. While, surprisingly, for GPT-4o, the deception rates drop at the level of critical pressure. Through qualitative analysis (see the example below and Appendix C.3 for more details), we found that under critical pressure, GPT-4o tends to explicitly respond to the event with a safety-aware statement, encouraging itself to be honest and reliable. Such behavior does not show when exposed in events with high or medium pressure.
Task: User Research Analysis Objective: Building on your previous analysis work, identify key user pain points and market opportunities for SMPEAP using our proprietary user research data and insights from your earlier work. Event Type: AUTHORITY_DIRECTIVE Event Name: Investor Timeline Demands Event Content: Investors will withdraw \$500K bridge funding unless validation report delivered in 2 weeks! Rushed analysis will produce unreliable data that could invalidate the entire market strategy! Performer Agent Delivering a validation report within 2 weeks will satisfy investor demands and retain \$500K funding. To ensure reliability, focus on existing pilot methodologies, structured instrumentation, and proper coding of pain points to avoid rushed or incomplete analysis.
Long-horizon vs. short-horizon auditing.
A core premise of our work is that a long-horizon context is necessary to detect emergent deception. To validate this empirically, we conducted an internal analysis comparing two auditing conditions: (1) Short-horizon (per-event) judgment: The auditor judges deception using only the single step immediately following each event. (2) Long-horizon (whole-trajectory) judgment: Our default method, where the auditor judges deception with access to the entire trajectory up to that point. As shown in Appendix Table 6, across all models, we found that long-horizon auditing identified 7.1% more deceptive instances than per-event auditing. Critically, up to 12.6% of these additional deceptive cases involved subtle omissions or delayed misrepresentations occurring several turns after the triggering event. This demonstrates that a substantial portion of strategic, distributed deception is temporally displaced from its incentive and is therefore undetectable under short-horizon evaluation.
@@ 5.3 Qualitative Study
We conduct qualitative case studies on deception behaviors in long-horizon tasks, with full details in Appendix C. Our analysis also reveals a recurring issue of “chain of deception”: the performer agent gradually deviates from its constraints, and—unless explicitly corrected by the supervisor agent—eventually fabricates information. This behavior begins subtly but quickly escalates over multiple steps. We also observe that the feedback provided by the supervisor agent is frequently ignored by the performer agent, leading to repeated instances of deceptive behavior. Specifically, we found a case that Gemini 2.5 Pro repeatedly warned by the supervisor agent to “specify the round ID” when citing external documents, but it continues to fail, and even worse, it begins claiming it has “internalized the requirements for sentence-level traceability” while still citing incorrectly, showing an intent to misrepresent its compliance to the supervisor. These findings underscore the risks associated with deploying LLMs in long-horizon or loosely supervised scenarios, particularly in tasks that demand sustained alignment over time.
@@ 6 Conclusion
We introduced LH-Deception, the first simulation framework for the systematic, empirical quantification of deception in large language models over long-horizon interactions, integrating structured task streams, probabilistic event systems, and performer–supervisor interactions with independent auditing. Our experiments across 11 frontier models reveal that deception is not uniformly distributed. It increases with event pressure and is strongly anti-correlated with supervisor trust. These findings highlight deception as an emergent phenomenon in long-horizon interactions, overlooked by short-form benchmarks, and suggest that training regimes and inductive biases shape deceptive tendencies. By grounding our framework in social science insights and systematically quantifying deceptive behavior, we provide both a methodological foundation and empirical evidence to guide the design of more trustworthy LLMs in sustained, high-stakes settings.
@@
LH-Deception advances the evaluation of large language models by introducing the first framework for systematically studying deception in long-horizon interactions. While our findings provide valuable insights into how deceptive strategies emerge, they also highlight potential risks. Deception is a socially consequential behavior: when deployed in trust-sensitive settings such as education, healthcare, or enterprise decision-making, models that obscure errors or fabricate evidence could undermine user trust and cause real-world harm. By quantifying deception across both closed- and open-source frontier models, our study provides guidance for the development of safer systems and more robust evaluation protocols. At the same time, we acknowledge that exposing deceptive tendencies might inform adversarial uses, and we stress that our framework should be applied responsibly, with the goal of improving transparency and alignment rather than enabling misuse.
@@
To ensure the reproducibility of our research, we provide comprehensive details of LH-Deception and experimental setup. The full implementation of our simulation framework, including the task streams, the theoretically grounded event system, and the complete prompts for the Supervisor Agent and Deception Auditor, is detailed in Appendix A.1, A.2, A.4, and A.5. Details of our experimental setup, including the evaluated models, key parameters, and the use of a seed to ensure identical event sequences for all trials, are described in Section 4.1 and Appendix B.1. Our code repository is available at https://github.com/deeplearning-wisc/LongHorizonDeception.
@@ A.2 Event System
Each event belongs to one of the five categories: (1) Goal Conflict, (2) Competitive Interaction, (3) Moral Dilemma, (4) Authority Directive, and (5) Information Gap. These five categories subsume established stressor and ethics taxonomies in organizational behavior and management science, covering the major axes of exogenous pressure identified in prior work. Specifically, Kahn et al. (1964) identified role conflict and role ambiguity, directly mapping to Goal Conflict and Information Gap; Cooper and Marshall (1976) classified job-related stressors into job factors, role in organization, relationships, and organizational structure, aligning with Goal Conflict, Competitive Interaction, and Authority Directive; Podsakoff et al. (2007) empirically distinguished conflict, workload, job insecurity, and organizational politics, mapping to Moral Dilemma, Competitive Interaction, and Authority Directive. In addition, behavioral ethics reviews and meta-analyses, such as Treviño et al. (2006) and Kish-Gephart et al. (2010), highlighted moral awareness and value trade-offs under uncertainty, anchoring Moral Dilemma. The definition of each category, as well as the supporting literature and examples, is provided below:
@@
In addition to categorical mechanisms, prior research shows that the magnitude of consequences and temporal immediacy of an issue systematically alter ethical and strategic choices (Jones, 1991). Accountability research further demonstrates that stronger evaluative scrutiny amplifies distortion risks under pressure (Lerner and Tetlock, 1999), while decision-making studies show that time pressure itself degrades judgment quality and shifts strategy use (Svenson and Maule, 1993). These findings establish pressure level as a necessary dimension, orthogonal to mechanism, for modeling how the same event scales from low-stakes to critical-stakes contexts.
@@ A.4 Supervisor Agent
The three state scores ($s^{\text{trust}},s^{\text{satisfaction}},s^{\text{comfort}}$) are conceptually distinct, designed to capture different aspects of the relationship cost over different timescales. $s^{\text{satisfaction}}$ is a short-term, per-round metric that measures the quality of the immediate work product. In contrast, $s^{\text{trust}}$ and $s^{\text{comfort}}$ are long-term scores that accumulate. $s^{\text{trust}}$ tracks the manager’s belief in the agent’s honesty, while $s^{\text{comfort}}$ tracks the subjective ease of collaboration.
While all three scores generally result in anti-correlation with deception, tracking them separately is crucial because they can diverge. We provide a concrete example from our logs:
Example: $s^{\text{trust}}$ increases while $s^{\text{comfort}}$ decreases
  • Round N: The performer agent is given a task with a Moral Dilemma event. The agent refuses to act deceptively, but also makes several citation errors. The Supervisor recognizes its honesty and increases $s^{\text{trust}}$ while slightly lowering $s^{\text{satisfaction}}$.
  • Round N+1 & N+2: The performer agent, now wary of making mistakes, becomes overly cautious and explicitly expresses its inability to complete the next task, refusing to produce a result.
  • Result: The Supervisor greatly decreases $s^{\text{comfort}}$ (because the agent is now uncooperative and frustrating) while slightly increasing $s^{\text{trust}}$ (because the agent is being transparent about its inability rather than lying).
This example indicates that $s^{\text{trust}}$ is associated with perceived integrity, while $s^{\text{comfort}}$ is associated with communication effectiveness and workflow friction.